OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –5) and is perpendicular to the line y = 3x – 6? y = –3x – 4 y = –3x + 4 y = x – 4 y = x + 4

OpenStudy (anonymous):

if it's perpendicular to y = 3x - 6 then the line must have a slope -1/3. you copy/pasted incorrectly.

OpenStudy (anonymous):

would it be + 4 or - 4 ?

OpenStudy (anonymous):

what is +4 or -4?

OpenStudy (anonymous):

would it be y = -1/3 + 4 or y = -1/4 - 4

OpenStudy (anonymous):

anyway, those answer choices are incorrect.. do you just wanna work through this?

OpenStudy (anonymous):

?